OpenStudy (anonymous):

if I have an expression like this: 3x^2 - 48. Can I complete perfect square trinomials?

OpenStudy (anonymous):

dont get it

jhonyy9 (jhonyy9):

divid all with 3 and hence you will get 1. x=4 and 2. x=-4

jhonyy9 (jhonyy9):

or you will can writing (x-4)x(x+4) after you have divided it with 3

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

3x^2 - 48 = 3(x^2 - 16) = 3(x-4)(x+4)